A specialty coffee store creates a special mix of two coffees at two different prices: $6.42 and $7.28 per pound each. How many pounds of the $7.28 coffee should be mixed with 7 pounds of $6.42 coffee to sell the mixture for $7.20 per pound?
Answer by sujatha.krishna(11)

Assume x is the number of pounds of 7.28 coffee.
So x pounds of 7.28 is mixed with 7 pounds of 6.42 coffee to get a per pound rate of 7.20. So the formula to calculate will be -
(x*7.28) + (7*6.42)
--------------------------- = 7.20
7+x
7.28x + 44.94 = 7.20 (7 +x)
7.28x + 44.94 = 50.4 + 7.20x
.08x = 5.46
x = 68.25